# Estimating Machine Translation Difficulty

### v1

@@ Appendix D Complete Results
Internal Artificial Crowd (XCOMET-QE-XXL) 3 0.221 0.203 0.271 0.194 0.213 0.273 0.173 0.238 0.195 0.233
Internal Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 4 0.207 0.211 0.256 0.184 0.212 0.221 0.176 0.229 0.175 0.203
Sentinel-MQM-24 5 0.182 0.167 0.216 0.169 0.173 0.220 0.142 0.204 0.150 0.197
External Artificial Crowd (XCOMET-QE-XXL) 6 0.177 0.175 0.192 0.146 0.179 0.240 0.128 0.194 0.160 0.183
Sentinel-MQM 6 0.175 0.164 0.205 0.159 0.171 0.223 0.118 0.201 0.141 0.190
External Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 7 0.166 0.181 0.174 0.121 0.162 0.247 0.128 0.180 0.136 0.162
@@
Internal Artificial Crowd (XCOMET-QE-XXL) 4 0.278 0.309 0.208 0.315
Internal Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 5 0.248 0.268 0.192 0.284
Sentinel-MQM-24 5 0.246 0.278 0.168 0.291
Sentinel-MQM 6 0.235 0.273 0.165 0.268
@@
External Artificial Crowd (XCOMET-QE-XXL) 8 0.207 0.243 0.159 0.220
External Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 9 0.185 0.209 0.145 0.201
@@
Sentinel-MQM-24 79.1 12.1%
External Artificial Crowd (XCOMET-QE-XXL) 78.3 13.3%
@@ Appendix E Oracles
Oracle methods use the true human judgments used to derive difficulty scores, as detailed in Section˜3:
  • “Oracle” assigns to each source text $x$ the true $d_{m,l}(x)$, for each $m$ and $l$.
  • “Oracle (source text + target lang)” estimates the difficulty of $x$ by average the true $d_{m,l}(x)$ across all translators ($\forall m\in\mathcal{M}_{l}$), meaning that its estimates do not vary across translators, but only across target languages.
  • “Oracle (source text only)” averages the true scores also across target languages, assigning the same score to each source text.
@@ Appendix F Creating Difficult Benchmarks – Quantitative Results
Results confirm the strong performance of our dedicated difficulty estimation model, Sentinel-MQM-24, which achieves substantially lower AvgScore and %Perfect values than random selection. It also achieves the best results among all automatic methods that rely solely on the source text. In particular, in Table D, it is outperformed in AvgScore only by External Artificial Crowd (XCOMET), a more computationally intensive approach that requires translating each source text with multiple large MT models and evaluating those translations using an XXL MT metric. Furthermore, External Artificial Crowd methods can produce difficulty scores conditioned on the target language, unlike Sentinel-MQM-24, which relies exclusively on the source text. On the other hand, Sentinel-MQM-24 obtains the best %Perfect score in Table D. In Table D, Sentinel-MQM-24 outperforms all automatic methods in both AvgScore and %Perfect, including External Artificial Crowd.
@@
Tables D and D provide a fine-grained breakdown of results across the WMT 24 domains (News, Social, Literary, and Speech) for the ESA and MQM test sets, respectively. These results show that the overall patterns hold consistently across domains. While absolute performance varies, Sentinel-MQM-24 and External Artificial Crowd achieve the strongest results in nearly all domain-specific evaluations.
This analysis supports the practical utility of difficulty estimation for controlled test set construction and confirms that learned estimators such as Sentinel-MQM-24 offer effective and reliable means for identifying source segments where MT systems are more likely to struggle.
@@ Appendix G Related work for Benchmark Creation
For machine translation specifically, Zhan et al. (2021a) use proxy of machine translation difficulty to inform better evaluation. Again, Zouhar et al. (2025a) automatically remove examples that are too easy from the evaluation set, corresponding to our internal artificial crowd with quality estimation.

### v2

@@ Appendix D Complete Results
True Crowd (XCOMET-QE-XXL) 3 0.221 0.203 0.271 0.194 0.213 0.273 0.173 0.238 0.195 0.233
True Crowd (MetricX-24-Hybrid-QE-XXL) 4 0.207 0.211 0.256 0.184 0.212 0.221 0.176 0.229 0.175 0.203
Sentinel-src-24 5 0.182 0.167 0.216 0.169 0.173 0.220 0.142 0.204 0.150 0.197
Artificial Crowd (XCOMET-QE-XXL) 6 0.177 0.175 0.192 0.146 0.179 0.240 0.128 0.194 0.160 0.183
Sentinel-src 6 0.175 0.164 0.205 0.159 0.171 0.223 0.118 0.201 0.141 0.190
Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 7 0.166 0.181 0.174 0.121 0.162 0.247 0.128 0.180 0.136 0.162
@@
True Crowd (XCOMET-QE-XXL) 4 0.278 0.309 0.208 0.315
True Crowd (MetricX-24-Hybrid-QE-XXL) 5 0.248 0.268 0.192 0.284
Sentinel-src-24 5 0.246 0.278 0.168 0.291
Sentinel-src 6 0.235 0.273 0.165 0.268
@@
Artificial Crowd (XCOMET-QE-XXL) 8 0.207 0.243 0.159 0.220
Artificial Crowd (MetricX-24-Hybrid-QE-XXL) 9 0.185 0.209 0.145 0.201
@@
Sentinel-src-24 79.1 12.1%
Artificial Crowd (XCOMET-QE-XXL) 78.3 13.3%
@@ Appendix E Oracles
Oracle methods use the true human judgments used to derive difficulty scores, as detailed in Section˜3. We consider three oracles that differ in the type of information they have access to:
  • Oracle (source text + target language + target translation) assigns to each source text $x$ the true $d_{m,l}(x)$, for each $m$ and $l$.
  • Oracle (source text + target language) estimates the difficulty of $x$ by averaging the true $d_{m,l}(x)$ across all models ($\forall m\in\mathcal{M}_{l}$), meaning that its estimates do not vary across models, but only across target languages.
  • Oracle (source text only) averages the true $d_{m,l}(x)$ across both models and target languages, assigning the same score to each source text regardless of target language or translator.
@@ Appendix F Creating Difficult Benchmarks – Quantitative Results
Results confirm the strong performance of our dedicated difficulty estimation model, Sentinel-src-24, which achieves substantially lower AvgScore and %Perfect values than random selection. It also achieves the best results among all automatic methods that rely solely on the source text. In particular, in Table D, it is outperformed in AvgScore only by Artificial Crowd (XCOMET), a more computationally intensive approach that requires translating each source text with multiple large models and evaluating those translations using an XXL MT metric. Furthermore, Artificial Crowd methods can produce difficulty scores conditioned on the target language, unlike Sentinel-src-24, which relies exclusively on the source text. On the other hand, Sentinel-src-24 obtains the best %Perfect score in Table D. In Table D, Sentinel-src-24 outperforms all automatic methods in both AvgScore and %Perfect, including Artificial Crowd.
@@
Tables D and D provide a fine-grained breakdown of results across the WMT 24 domains (News, Social, Literary, and Speech) for the ESA and MQM test sets, respectively. These results show that the overall patterns hold consistently across domains. While absolute performance varies, Sentinel-src-24 and Artificial Crowd achieve the strongest results in nearly all domain-specific evaluations.
This analysis supports the practical utility of difficulty estimation for controlled test set construction and confirms that learned estimators such as Sentinel-src-24 offer effective and reliable means for identifying source segments where MT systems are more likely to struggle.
@@ Appendix G Related work for Benchmark Creation
For machine translation specifically, Zhan et al. (2021a) use proxy of machine translation difficulty to inform better evaluation. Again, Zouhar et al. (2025a) automatically remove examples that are too easy from the evaluation set, corresponding to our True Crowd with quality estimation.